[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Sat Oct 6 11:20:02 EDT 2018

Logentry Text:
During the controlled access yesterday the following was accomplished:

--> SHMS aerogel tray swap [https://logbooks.jlab.org/entry/3604756] in record time! We were ready to go to beam permit by 1 pm; the magic was delivered by Vlad, Jerry, Stan, Larry, Rob

--> thanks to Howard most ringing in HMS drift chambers got fixed [https://logbooks.jlab.org/entry/3604763]

--> the HMS dipole JT valve that's regulating the LN2 level got fixed so we don't have to regulate manually 

--> Hall C dump diffuser investigation [https://logbooks.jlab.org/entry/3604658]

--> Dave Meekins looked into what might cause the target ice ball formation

During the access the momenta and angles of the spectrometers were changed so we can take data at the next physics kinematic setting (Q2=4.4 W=2.74 x=0.4 ).  We started the first run on LH2 at the new kinematic setting at 13:24 pm. Rates are such at these kinematics that we can take the highest current that Ops can send us. A history of the beam current delivered since yesterday is shown here [https://logbooks.jlab.org/entry/3605108]. 

There were several major interruptions in beam delivery. One happened yesterday shortly after we got beam back following the controlled access. We lost beam at about 14:00 pm due to yet another PSS drop. We got beam back around 16:00 and we resumed production and we collected about 58% of the min statistical goal on LH2 at  Q2=4.4 W=2.74 x=0.4 theta_Shms = 9.81 deg by 2:20 am when beam went away for almost 9 hours.
 An ion pump faulted closing 2 valves around 2L23 (VBV2L23A and VBV2L23B). Eventually experts came on site and tracked the problem to be with a bad HV cable on that ion pump. The cable has been replaced this morning  and we are now waiting for beam.

We will be moving to the next SHMS angle as soon as we take the Al data at the current setting. We are trying to roughly stick to 2 shifts per SHMS angle at this setting when possible.

A link to the data taking progress can be found here [https://hallcweb.jlab.org/wiki/index.php/Kaon_LT_Status_Page]
There will be no RC meeting today.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3605118
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181006/cdd09fcb/attachment.html>

More information about the Hallc_running mailing list